OpenStudy (jazzyluvsu123):

Sally and her husband Jim left their house at the same time traveling in the same direction. Sally drove at 55 MPH and Jim drove at 65 mph. How many hours was it before Jim was 100 miles ahead of Sally?

OpenStudy (daniellelovee):

100/55 100/65 since they are miles per hour

OpenStudy (daniellelovee):

their equations are 55h=100 and 65h=100 therefore you hve to use PEMDAS backwards to find h and then subtract

OpenStudy (daniellelovee):

do that first and then we'll do the last step

OpenStudy (jazzyluvsu123):

So the whole equation would be 55x=100-65x=100?

OpenStudy (daniellelovee):

65x=100-55x=100 since we want to know how ahead was the jim

OpenStudy (jazzyluvsu123):

When i did 65x=100 i got 20/13??

OpenStudy (jazzyluvsu123):

then 20/11 for the other

OpenStudy (daniellelovee):

exactly so just subtract them

OpenStudy (jazzyluvsu123):

-40/143???

OpenStudy (daniellelovee):

that's an odd number hold on

OpenStudy (daniellelovee):

do you have options?

OpenStudy (jazzyluvsu123):

No i have to type the answer in. I need to make an equation to answer it but i can't seem to get the right equation

OpenStudy (jazzyluvsu123):

I was gonna do 55x=65(x *then i got stuck*

OpenStudy (jazzyluvsu123):

Do i do x-100?

OpenStudy (daniellelovee):

yeah I think we would have to subtract her-husband to get a positive number

OpenStudy (jazzyluvsu123):

i don't think 650 is the answer? or do i change that into hours maybe that is in minutes

OpenStudy (daniellelovee):

I don't know I think the answer would be 0.28 but I'm unsure

OpenStudy (jazzyluvsu123):

650/60 is 10.8333 Repeating

OpenStudy (daniellelovee):

ohhh wait true is that it says until he is ahead by 100 miles

OpenStudy (jazzyluvsu123):

yes

OpenStudy (dumbcow):

distance = rate*time 65x - 55x = 100 10x = 100 x = 10

OpenStudy (jazzyluvsu123):

Ooo lemme try that equation

OpenStudy (daniellelovee):

55x20=1100 65x20=1300 so that would be 200 miles ahead

OpenStudy (daniellelovee):

yup it is 10

OpenStudy (daniellelovee):

since 55x10=550 and 65x10=650 and 650-550=100

OpenStudy (jazzyluvsu123):

@dumbcow *even tho ur not dumb ..* thanks so much that worked and thank you @Daniellelovee for your help swell (:

OpenStudy (daniellelovee):

yeah lol no problem I just confused you xD